# appendix,by

### v1

@@ 3 Amalgamation
An amalgamation problem of (enriched) valued fields induces amalgamation problems at the residue field and value group levels. Therefore, an obvious obstruction to the solution of an amalgamation problem is that one of these induced problems cannot be solved.
In the case of valued difference fields, value difference groups can always be amalgamated by [DM24, Theorem 5.10], but residual obstructions may still be present.
@@
Amalgamation problems may fail to have solutions even in the absence of obstacles coming from the residue field (e.g. if the latter is algebraically closed, see ˜3.28 and the following discussion).
@@
Let $k$ be algebraically closed, let $\sigma\in\operatorname{Aut}(k)$ be arbitrary, and consider the Hahn product $k((\mathbb{R}))$ of $(k,\sigma)$ and $(\mathbb{R},0,+,\leq,\mathrm{id}_{\mathbb{R}})$. Let $K_{0}\coloneqq k((\mathbb{Z}))\subseteq k((\mathbb{R}))$, and let $K\coloneqq K_{0}(t^{\pi})$. By the Zariski–Abhyankar inequality [Dri14, Corollary 3.25], the residue field of $K$ is an algebraic extension of $k$, so it is equal to $k$. Let $L$ and $M$ be valued fields isomorphic to $K(\sqrt{t^{\pi}})$, endowed with automorphisms $\sigma_{L}$ and $\sigma_{M}$ extending the automorphism on $K$ so that $\sigma_{L}$ fixes the square roots of $t^{\pi}$ while $\sigma_{M}$ swaps them. Then $(L,\sigma_{L})$ and $(M,\sigma_{M})$ cannot be amalgamated over $K$.
This is not surprising, as similar issues are well known to exist when one tries to amalgamate e.g. henselian valued fields of residual characteristic 0 (with no extra structure). Common solutions are to expand the language by an angular component or a cross-section, which in our case will be assumed to be equivariant. In ˜3.28 we will show that, in the residual characteristic 0 case, this allows to characterise which amalgamation problems have solutions. Our result also works when a lift of the residue field is added to the language.
@@ 3.1 Set-up
A difference field $(k,\sigma)$ is linearly difference closed iff for all $n\in\mathbb{N}$ and $\alpha_{0},\ldots,\alpha_{n}\in k$ with $\alpha_{n}\neq 0$ the equation $1+\alpha_{0}x+\ldots+\alpha_{n}\sigma^{n}(x)$ has a solution in $k$.
@@
The following definitions appear for example in [AD11, Section 2].
@@
Let $(K,\sigma)$ be a difference field, $(K_{0},\sigma_{|K_{0}})$ a difference subfield. An element $a\in K$ is $\sigma$-algebraic over $K_{0}$ iff there is a nonzero $\sigma$-polynomial $G$ over $K_{0}$ such that $G(a)=0$, and $\sigma$-transcendental otherwise. If $a$ is $\sigma$-algebraic over $K_{0}$, a minimal $\sigma$-polynomial for $a$ over $K_{0}$ is a nonzero $\sigma$-polynomial $G$ of minimal complexity such that $G(a)=0$.
For $G$ a $\sigma$-polynomial, we refer to [AD11, Section 2] for the definition of the $\sigma$-polynomials $G_{(\mathbf{i})}$. Note that if $\mathbf{i}\neq 0$ then $G_{(\mathbf{i})}$ has strictly lower complexity than $G$.
For $G$ a $\sigma$-polynomial and $a\in K$, we refer to [Azg10, Definition 4.2] or [DO15, Definition 4.1] for the definitions of $(G,a)$ being in $\sigma$-Hensel configuration and of the value $\gamma(G,a)$. Recall that a valued difference field is $\sigma$-henselian iff, whenever $(G,a)$ is in $\sigma$-Hensel configuration, there is $b$ with $G(b)=0$ and $v(a-b)=\gamma(G,a)$. [^11] There is another notion of $\sigma$-henselianity, designed to work for valued difference fields with analytic structure, see [Rid17, Definition 4.10]. In the case where the analytic structure is trivial and $\operatorname{char}(k)=0$, the two notions are equivalent, as they both imply (and are implied by) elementary equivalence to a maximally complete valued difference field with linearly difference closed residue difference field. We thank Silvain Rideau-Kikuchi for pointing this out. Our uses of $\sigma$-henselianity will be mediated by the following observation.
Fact 3.6 ([DO15, Remark 4.2]).
@@
This is [AD11, Theorem 4.8]. Note that its proof does not assume that the automorphism $\sigma$ is isometric (i.e. $\sigma(\gamma)=\gamma$ for all $\gamma\in\Gamma$), even if the main results of the cited paper do. Moreover, even though the notion of $\sigma$-henselianity used there is different, it can be seen by using ˜3.6 that the proof still works with the notion of $\sigma$-henselianity used here. ∎
@@
Let $\mathcal{L}$ be either $\mathcal{L}_{\mathrm{\sigma}}$, $\mathcal{L}_{\operatorname{ac},\sigma}$, or the expansion of one of these by a lift. Every existentially closed $\mathcal{L}$-valued difference field of residual characteristic 0 is $\sigma$-henselian.
@@
If $\mathcal{L}$ equals $\mathcal{L}_{\mathrm{\sigma}}$ (resp., $\mathcal{L}_{\operatorname{ac},\sigma}$), this follows from Proposition˜2.3(3) (resp., Corollary˜2.22), the fact that every $\sigma$-algebraically maximal (i.e. having no proper $\sigma$-algebraic immediate extension) valued difference field with linearly difference closed residue field and residual characteristic 0 is $\sigma$-henselian [DO15, Corollary 5.6(2)], and the fact that $\sigma$-henselianity is expressible by a set of $\forall\exists$-sentences in $\mathcal{L}$. The latter can be checked by inspection, by observing that, in [DO15, Definition 4.1], for $G$ of fixed complexity, it suffices to check finitely many of the $G_{(\mathbf{j})}$ because the other ones will be null regardless of the coefficients of $G$.
@@
    If $(K,v,\sigma,s,\iota)$ is existentially closed, and $\operatorname{ac}(x)=\operatorname{res}(x/s(v(x)))$, then $(K,v,\sigma,\operatorname{ac},\iota)$ is existentially closed.
@@
For $\mathcal{L}_{\operatorname{ac},\sigma}$ we use that, by ˜3.8, partial lifts of the residue field may always be extended. This also implies the result for $\mathcal{L}_{\sigma}$ by Corollary˜2.19, and for $\mathcal{L}_{\operatorname{ac}}$ and $\mathcal{L}$ by Corollary˜2.22. For $\mathcal{L}_{\operatorname{ac},\iota}$ we apply [Kes24, Proposition 1.8], which allows us to repeat the argument with quantifier elimination as in Proposition˜2.3.
@@
Let $\mathcal{L}$ be $\mathcal{L}_{\sigma},\mathcal{L}_{\operatorname{ac},\sigma},\mathcal{L}_{s,\sigma}$, or an expansion of the previous by a lift. If $K$ is an existentially closed $\mathcal{L}$-valued difference field of residual characteristic 0, then it is $\sigma$-henselian.
@@
The following facts about valued difference fields appear in Sections 2 and 6 of [AD11]. As in ˜3.8, the isometricity assumption is not used in these three lemmas, hence they hold for arbitrary valued difference fields.
@@
The proof of [AD11, Lemma 6.3] does not use the elementarity requirements in the definition of a good map in [AD11, Section 6], thus we can relax the assumption of being a good map to being an isomorphism of $\mathrm{ac}$-valued fields. ∎
@@
This is essentially [DO15, Lemma 3.2], together with the observation that v(ca)>v(a)v(c-a)>v(a) if and only if $(\operatorname{ac}(a),v(a))=(\operatorname{ac}(c),v(c))$. The latter can be easily shown by using that the pair of maps $(\operatorname{ac},v)$ induces an isomorphism of $k\times\Gamma$ with the $\mathrm{RV}$-structure $\mathrm{VF}^{\times}/(1+\mathfrak{m})$, where $\mathfrak{m}$ denotes the maximal ideal. ∎
@@ 3.2 Amalgamation
From now on, all fields will be assumed to have characteristic 0.
@@
This is a variant of [Pal12, Lemma 8.8]. One modifies its proof by replacing the tacitly used [Pal12, Lemma 8.5] by [DO15, Lemma 3.6] (as in [DO15, Lemma 6.1] [^12] Point (ii) of the cited lemma incorrectly states that, in its notation, the residue field of $E\langle a\rangle$ is the same as that of $E$. In fact, the residue field will in general grow. We thank Simone Ramello for bringing this fact to our attention.) and observes that the saturation assumption in [Pal12, Lemma 8.8] is only used to realise a quantifier-free type in $U$. ∎
@@
We now show how to extend $k$-elementary maps to make sure that the valuation and residue maps on the domain substructure are surjective (note that, in the absence of a cross-section, the valuation map in a substructure need not be surjective, and similarly for lifts and residue).
@@
Suppose now $\mathcal{L}=\mathcal{L}_{s,\iota,\sigma}$. For all $i\in\mathbb{Z}$, let $a_{i}\coloneqq\iota(\sigma^{i}(\alpha))$ and $b_{i}\coloneqq\iota(\sigma^{i}(\beta))$. We prove by induction on $i\in\mathbb{N}$ that the $k$ sort of the $\mathcal{L}_{s,\iota}$-structure generated by $k(A)\cup\operatorname{VF}(A)(a_{-i},\ldots,a_{i})$ equals $k(A)(a_{-i},\ldots,a_{i})$ and that there is an $\mathcal{L}_{s,\iota}$-isomorphism $f_{i}$ extending $f$ and sending the tuple $(a_{-i},\ldots,a_{i})$ to $(b_{-i},\ldots,b_{i})$. Suppose this holds for some $i\in\mathbb{N}$. Applying [Kes24, Proposition 1.7] we obtain an $\mathcal{L}_{s,\iota}$-isomorphism extending $f_{i}$ and sending $a_{i+1}$ to $b_{i+1}$. Applying [Kes24, Proposition 1.7] again, this time to $a_{-i-1}$, we obtain the desired $k$-elementary map $f_{i+1}$. We now conclude by taking $f^{\prime}\coloneqq\bigcup_{i\in\mathbb{N}}f_{i}$.
Suppose now that $\mathcal{L}=\mathcal{L}_{\operatorname{ac},\sigma}$. We may assume $\alpha\in k(A)$ and $\beta\in k(B)$. If $\alpha$ is $\sigma$-transcendental over $\operatorname{res}(\mathcal{O}(A))$, then $\beta$ is $\sigma$-transcendental over $\operatorname{res}(\mathcal{O}(B))$. Pick $a\in\mathcal{O}(U)$ and $b\in\mathcal{O}(N)$ with $\operatorname{res}(a)=\alpha$ and $\operatorname{res}(b)=\beta$. By ˜3.14(3) $f$ extends to an $\mathcal{L}_{\sigma}$-isomorphism $f^{\prime}\colon A\langle a\rangle\to B\langle b\rangle$ sending $a$ to $b$, and by ˜3.14(2) $\Gamma(A)=\Gamma(A\langle a\rangle)$. By the choice of $\beta$, $f^{\prime}|_{k(A)}$ is an elementary map $k(U)\to k(N)$. We then use ˜3.16 to see that $f^{\prime}$ commutes with $\operatorname{ac}$. Thus, $f^{\prime}$ is a $k$-elementary map.
@@
If $\mathcal{L}=\mathcal{L}_{s,\sigma}$, we use the same proof as in the previous case, except that we do not need to invoke ˜3.16, as commutation with the cross-section is given by the fact that $\Gamma$ does not grow. ∎
@@
Suppose $\mathcal{L}=\mathcal{L}_{\operatorname{ac},\sigma}$. By ˜3.7 we are in the assumptions of Lemma˜3.20, hence there is $a\in U$ regular over $\operatorname{VF}(A)$ with $v(a)=\gamma$. By Lemma˜3.24, we may extend $f$ to some element of $\mathcal{O}(U)$ with residue $\operatorname{ac}(a)$. Pick $b\in N$ with $v(b)=\delta$ and $\operatorname{ac}(b)=f(\operatorname{ac}(a))$. By ˜3.18, $b$ is regular over $B$. By Lemma˜3.20, we may extend $f$ to a field isomorphism sending $a$ to $b$. From the fact that $\operatorname{ac}(b)=f(\operatorname{ac}(a))$ and Lemma˜3.21, we conclude that $f$ commutes with $\operatorname{ac}$.
If $\mathcal{L}=\mathcal{L}_{\operatorname{ac},\iota,\sigma}$, argue as in the previous case and then note that $f$ then also commutes with the lift by regularity of $a$.
@@
If instead there is some positive $\ell\in\mathbb{N}$ such that $\ell\sigma^{i+1}(\gamma)\in\Gamma(A)+\mathbb{Z}\sigma^{-i}(\gamma)+\ldots+\mathbb{Z}\sigma^{i}(\gamma)$, and take the minimal such $\ell$, together with integer coefficients $h_{i}$, and $\gamma_{0}\in\Gamma(A)$ such that $\ell\sigma^{i+1}(\gamma)=\gamma_{0}+\sum_{-i\leq j\leq i}h_{i}\sigma^{i}(\gamma)$. Applying the cross-section, we get that $a_{i+1}^{\ell}=s(\gamma_{0})\prod_{-i\leq j\leq i}a_{j}^{h_{j}}$ and $b_{i+1}^{\ell}=s(f(\gamma_{0}))\prod_{-i\leq j\leq i}b_{j}^{h_{j}}$. It follows then by ˜3.13(1) that $x^{\ell}-s(\gamma_{0})\prod_{-i\leq j\leq i}a_{j}^{h_{j}}$ is the minimal polynomial of $a_{i+1}$ over $\operatorname{VF}(A\langle a_{-i},\ldots,a_{i}\rangle)$, and $x^{\ell}-f(s(\gamma_{0}))\prod_{-i\leq j\leq i}b_{j}^{h_{j}}$ is the minimal polynomial of $b_{i+1}$ over $\operatorname{VF}(B\langle b_{-i},\ldots,b_{i}\rangle)$. Because we have $f_{i}(\prod_{-i\leq j\leq i}a_{j}^{h_{j}})=\prod_{-i\leq j\leq i}b_{j}^{h_{j}}$, we get that $f_{i}$ extends to map
@@
If $\mathcal{L}=\mathcal{L}_{s,\iota,\sigma}$, the proof is analogous to the proof of the case $\mathcal{L}_{s,\iota,\sigma}$ in Lemma˜3.24 with the roles of the sorts $\Gamma$ and $k$ interchanged, the role of $\iota$ now played by $s$, and the applications of [Kes24, Proposition 1.7] replaced by applications of [Kes24, Proposition 2.1]. ∎
@@
Let $E,M,N$ be substructures of $\mathcal{L}$-valued difference fields of residual characteristic 0, with $M$ and $N$ extending $E$ and such that:
@@
Since $k(M)$ embeds into $k(N)$ over $k(E)$, there is an existentially closed (in particular, $\sigma$-henselian by Corollary˜3.11) and quantifier-free $\lvert M\rvert^{+}-$saturated extension $U$ of $M$ such that the restrictions to the algebraic closure of $k(E)$ of the automorphisms of $k(U)$ and $k(N)$ are isomorphic. By [CH99, Theorem 1.3] $\mathrm{id}_{E}$ is then a $k$-elementary map between substructures of $U$ and $N$.
@@
Observe that the proof of [DO15, Corollary 5.10] [^15] For the reader’s convenience, we point out that most of the proofs in [DO15, Section 5], which culminates in the corollary of interest, are omitted as they heavily rely on similar material presented in [AD11, Section 5] only uses saturation to realise sets of quantifier-free formulas. Namely, the only step where saturation is used is in the proof of [AD11, Lemma 5.9], in saying that, in our notation, certain pseudo-Cauchy sequences from $B_{1}$ have pseudolimits (see for example [DO15, Section 2] for the definitions of pseudo-Cauchy sequence, or pc-sequence, and of pseudolimit). To conclude, observe that if $\pi(x)$ is a set of quantifier-free formulas over $B_{1}$ stating that $x$ is a pseudolimit of a certain pc-sequence then, by definition of pc-sequence, $\pi(x)$ is finitely satisfiable in $B_{1}$, and a fortiori in $N$. ∎
As $k(A_{1})=k(A_{2})$, in order to conclude that $f_{2}$ is a $k$-elementary map, it remains to be shown that it commutes with $\operatorname{ac}$. This follows from the fact, because $a\in A_{2}$ is a pseudolimit of a pseudo-Cauchy sequence, there is $b\in A_{1}$ such that v(ba)>v(a)v(b-a)>v(a).
@@
An amalgamation problem of valued difference fields of residual characteristic 0 with either an angular component or a cross-section, and possibly with a lift has a solution if and only if the induced problem on the residue difference fields has a solution.

### v2

@@ 3 Amalgamation
An amalgamation problem of (expanded) valued fields induces amalgamation problems at the residue field and value group levels. Therefore, an obvious obstruction to the solution of an amalgamation problem is that one of these induced problems cannot be solved.
In the case of valued difference fields, value difference groups can always be amalgamated by [DM, Theorem 5.10], but residual obstructions may still be present.
@@
Even in the absence of obstacles coming from the residue field (e.g. algebraically closed difference fields are amalgamation bases, see ˜3.28 and the following discussion), amalgamation problems may fail to have solutions.
@@
Let $k$ be algebraically closed, let $\sigma\in\operatorname{Aut}(k)$ be arbitrary, and consider the Hahn product $k((\mathbb{R}))$ of $(k,\sigma)$ and $(\mathbb{R},0,+,\leq,\mathrm{id}_{\mathbb{R}})$. Let $K_{0}\coloneqq k((\mathbb{Z}))\subseteq k((\mathbb{R}))$, and let $K\coloneqq K_{0}(t^{\pi})$. By the Zariski–Abhyankar inequality [Dri, Corollary 3.25], the residue field of $K$ is an algebraic extension of $k$, so it is equal to $k$. Let $L$ and $M$ be valued fields isomorphic to $K(\sqrt{t^{\pi}})$, endowed with automorphisms $\sigma_{L}$ and $\sigma_{M}$ extending the automorphism on $K$ so that $\sigma_{L}$ fixes the square roots of $t^{\pi}$ while $\sigma_{M}$ swaps them. Then $(L,\sigma_{L})$ and $(M,\sigma_{M})$ cannot be amalgamated over $K$.
This is not surprising, as similar issues are well known to exist when one tries to amalgamate e.g. henselian valued fields of residual characteristic zero (with no extra structure). Common solutions are to expand the language by an angular component or a cross-section, which in our case will be assumed to be equivariant. In ˜3.28 we will show that, in the residual characteristic zero case, this allows to characterise which amalgamation problems have solutions. Our result also works when a lift of the residue field is added to the language.
@@ 3.1 Set-up
A difference field $(k,\sigma)$ is linearly difference closed iff for all $n\in\mathbb{N}$ and $\alpha_{0},\ldots,\alpha_{n}\in k$ with $\alpha_{n}\neq 0$ the equation $1+\alpha_{0}x+\ldots+\alpha_{n}\sigma^{n}(x)=0$ has a solution in $k$.
@@
The following definitions appear for example in [AvD, Section 2].
@@
Let $(K,\sigma)$ be a difference field, $(K_{0},\sigma_{|K_{0}})$ a difference subfield. An element $a\in K$ is $\sigma$-algebraic over $K_{0}$ iff there is a nonzero $\sigma$-polynomial $G$ over $K_{0}$ such that $G(a)=0$, and $\sigma$-transcendental over $K_{0}$ otherwise. If $a$ is $\sigma$-algebraic over $K_{0}$, a minimal $\sigma$-polynomial for $a$ over $K_{0}$ is a nonzero $\sigma$-polynomial $G$ of minimal complexity such that $G(a)=0$.
For $G$ a $\sigma$-polynomial, we refer to [AvD, Section 2] for the definition of the $\sigma$-polynomials $G_{(\mathbf{i})}$. Note that if $\mathbf{i}\neq 0$ then $G_{(\mathbf{i})}$ has strictly lower complexity than $G$.
For $G$ a $\sigma$-polynomial and $a\in K$, we refer to [Az, Definition 4.2] or [DO, Definition 4.1] for the definitions of $(G,a)$ being in $\sigma$-Hensel configuration and of the value $\gamma(G,a)$. Recall that a valued difference field is $\sigma$-henselian iff, whenever $(G,a)$ is in $\sigma$-Hensel configuration, there is $b$ with $G(b)=0$ and $v(a-b)=\gamma(G,a)$. [^10] There is another notion of $\sigma$-henselianity, designed to work for valued difference fields with analytic structure, see [Ri, Definition 4.10]. In the case where the analytic structure is trivial and $\operatorname{char}(k)=0$, the two notions are equivalent, as they both imply (and are implied by) elementary equivalence to a maximally complete valued difference field with linearly difference closed residue difference field. We thank Silvain Rideau-Kikuchi for pointing this out. Our uses of $\sigma$-henselianity will be mediated by the following observation.
Fact 3.6 ([DO, Remark 4.2]).
@@
This is [AvD, Theorem 4.8]. Note that its proof does not assume that the automorphism $\sigma$ is isometric (i.e. $\sigma(\gamma)=\gamma$ for all $\gamma\in\Gamma$), even if the main results of the cited paper do. Moreover, even though the notion of $\sigma$-henselianity used there is different, it can be seen by using ˜3.6 that the proof still works with the notion of $\sigma$-henselianity used here. ∎
@@
Let $\mathcal{L}$ be either $\mathcal{L}_{\mathrm{\sigma}}$, $\mathcal{L}_{\operatorname{ac},\sigma}$, or the expansion of one of these by a lift. Every existentially closed $\mathcal{L}$-valued difference field of residual characteristic zero is $\sigma$-henselian.
@@
If $\mathcal{L}$ equals $\mathcal{L}_{\mathrm{\sigma}}$ (resp. $\mathcal{L}_{\operatorname{ac},\sigma}$), this follows from Proposition˜2.3(3) (resp. Corollary˜2.22), the fact that every $\sigma$-algebraically maximal (i.e. having no proper $\sigma$-algebraic immediate extension) valued difference field with linearly difference closed residue field and residual characteristic zero is $\sigma$-henselian [DO, Corollary 5.6(2)], and the fact that $\sigma$-henselianity is expressible by a set of $\forall\exists$-sentences in $\mathcal{L}$. The last fact can be checked by inspection, by observing that, in [DO, Definition 4.1], for $G$ of fixed complexity, it suffices to check finitely many of the $G_{(\mathbf{j})}$ because the other ones will be null regardless of the coefficients of $G$.
@@
    If $(K,v,\sigma,s,\iota)$ is existentially closed, and $\operatorname{ac}(x)=\operatorname{res}(x/s(v(x)))$, then so is $(K,v,\sigma,\operatorname{ac},\iota)$.
@@
For $\mathcal{L}_{\operatorname{ac},\sigma}$ we use that, by ˜3.8, partial lifts of the residue field may always be extended. This also implies the result for $\mathcal{L}_{\sigma}$ by Corollary˜2.19, and for $\mathcal{L}_{\operatorname{ac}}$ and $\mathcal{L}$ by Corollary˜2.22. For $\mathcal{L}_{\operatorname{ac},\iota}$ we apply [Kesting, Proposition 1.8], which allows us to repeat the argument with quantifier elimination as in Proposition˜2.3.
@@
Let $\mathcal{L}$ be $\mathcal{L}_{\sigma},\mathcal{L}_{\operatorname{ac},\sigma},\mathcal{L}_{s,\sigma}$, or an expansion of the previous by a lift. If $K$ is an existentially closed $\mathcal{L}$-valued difference field of residual characteristic zero, then it is $\sigma$-henselian.
@@
The following facts about valued difference fields appear in Sections 2 and 6 of [AvD]. As in ˜3.8, the isometricity assumption is not used in these three lemmas, hence they hold for arbitrary valued difference fields.
@@
The proof of [AvD, Lemma 6.3] does not use the elementarity requirements in the definition of a good map in [AvD, Section 6], thus we can relax the assumption of being a good map to being an isomorphism of $\mathrm{ac}$-valued fields. ∎
@@
This is essentially [DO, Lemma 3.2], together with the observation that v(ca)>v(a)v(c-a)>v(a) if and only if $(\operatorname{ac}(a),v(a))=(\operatorname{ac}(c),v(c))$. This equality can be easily shown by using that the pair of maps $(\operatorname{ac},v)$ induces an isomorphism of $k\times\Gamma$ with the $\mathrm{RV}$-structure $\mathrm{VF}^{\times}/(1+\mathfrak{m})$, where $\mathfrak{m}$ denotes the maximal ideal. ∎
@@ 3.2 Amalgamation
From now on, all fields will be assumed to have characteristic zero.
@@
This is a variant of [Pa, Lemma 8.8]. One modifies its proof by replacing the tacitly used [Pa, Lemma 8.5] by [DO, Lemma 3.6] (as in [DO, Lemma 6.1] [^11] Point (ii) of the cited lemma incorrectly states that, in its notation, the residue field of $E\langle a\rangle$ is the same as that of $E$. In fact, the residue field will in general grow. We thank Simone Ramello for bringing this fact to our attention.) and observes that the saturation assumption in [Pa, Lemma 8.8] is only used to realise a quantifier-free type in $U$. ∎
@@
We now show how to extend $k$-elementary maps to make sure that the valuation and residue maps on the domain substructure are surjective (note that, in the absence of a cross-section, the valuation map in a substructure need not be surjective, and similarly for lift and residue).
@@
Suppose now $\mathcal{L}=\mathcal{L}_{s,\iota,\sigma}$. For all $i\in\mathbb{Z}$, let $a_{i}\coloneqq\iota(\sigma^{i}(\alpha))$ and $b_{i}\coloneqq\iota(\sigma^{i}(\beta))$. We prove by induction on $i\in\mathbb{N}$ that the $k$ sort of the $\mathcal{L}_{s,\iota}$-structure generated by $k(A)\cup\operatorname{VF}(A)(a_{-i},\ldots,a_{i})$ equals $k(A)(\sigma^{-i}(\alpha),\ldots,\sigma^{i}(\alpha))$ and that there is an $\mathcal{L}_{s,\iota}$-isomorphism $f_{i}$ extending $f$ and sending the tuple $(a_{-i},\ldots,a_{i})$ to $(b_{-i},\ldots,b_{i})$. Suppose this holds for some $i\in\mathbb{N}$. Applying [Kesting, Proposition 1.7] we obtain an $\mathcal{L}_{s,\iota}$-isomorphism extending $f_{i}$ and sending $a_{i+1}$ to $b_{i+1}$. Applying [Kesting, Proposition 1.7] again, this time to $a_{-i-1}$, we obtain the desired $k$-elementary map $f_{i+1}$. We now conclude by taking $f^{\prime}\coloneqq\bigcup_{i\in\mathbb{N}}f_{i}$.
Suppose now that $\mathcal{L}=\mathcal{L}_{\operatorname{ac},\sigma}$. We may assume $\alpha\in k(A)$ and $\beta\in k(B)$. If $\alpha$ is $\sigma$-transcendental over $\operatorname{res}(\mathcal{O}(A))$, then $\beta$ is $\sigma$-transcendental over $\operatorname{res}(\mathcal{O}(B))$. Pick $a\in\mathcal{O}(U)$ and $b\in\mathcal{O}(N)$ with $\operatorname{res}(a)=\alpha$ and $\operatorname{res}(b)=\beta$. By ˜3.14(3) $f$ extends to an $\mathcal{L}_{\sigma}$-isomorphism $f^{\prime}\colon A\langle a\rangle\to B\langle b\rangle$ sending $a$ to $b$, and by ˜3.14(2) $\Gamma(A)=\Gamma(A\langle a\rangle)$. By the choice of $\beta$, $f^{\prime}|_{k(A\langle a\rangle)}$ is an elementary map $k(U)\to k(N)$. We then use ˜3.16 to see that $f^{\prime}$ commutes with $\operatorname{ac}$. Thus, $f^{\prime}$ is a $k$-elementary map.
@@
If $\mathcal{L}=\mathcal{L}_{s,\sigma}$, we use the same proof as in the previous case, except that we do not need to invoke ˜3.16. Commutation with the cross-section is given by the fact that $\Gamma$ does not grow. ∎
@@
Suppose $\mathcal{L}=\mathcal{L}_{\operatorname{ac},\sigma}$. By ˜3.7 we are in the assumptions of Lemma˜3.20, hence there is $a\in U$ regular over $\operatorname{VF}(A)$ with $v(a)=\gamma$. By Lemma˜3.24, we may extend $f$ to some element of $\mathcal{O}(U)$ with residue $\operatorname{ac}(a)$. Pick $b\in N$ with $v(b)=\delta$ and $\operatorname{ac}(b)=f(\operatorname{ac}(a))$. By ˜3.18, $b$ is regular over $B$. By Lemma˜3.20, we may extend $f$ to a valued difference field isomorphism sending $a$ to $b$. From the fact that $\operatorname{ac}(b)=f(\operatorname{ac}(a))$ and Lemma˜3.21, we conclude that $f$ commutes with $\operatorname{ac}$.
If $\mathcal{L}=\mathcal{L}_{\operatorname{ac},\iota,\sigma}$, argue as in the previous case and then, to check that $f$ also commutes with the lift, use the fact $\operatorname{ac}(a)$ is transcendental by regularity of $a$.
@@
If instead there is some positive $\ell\in\mathbb{N}$ such that $\ell\sigma^{i+1}(\gamma)\in\Gamma(A)+\mathbb{Z}\sigma^{-i}(\gamma)+\ldots+\mathbb{Z}\sigma^{i}(\gamma)$, take the minimal such $\ell$, together with integer coefficients $h_{i}$, and $\gamma_{0}\in\Gamma(A)$ such that $\ell\sigma^{i+1}(\gamma)=\gamma_{0}+\sum_{-i\leq j\leq i}h_{i}\sigma^{i}(\gamma)$. Applying the cross-section, we get that $a_{i+1}^{\ell}=s(\gamma_{0})\prod_{-i\leq j\leq i}a_{j}^{h_{j}}$ and $b_{i+1}^{\ell}=s(f(\gamma_{0}))\prod_{-i\leq j\leq i}b_{j}^{h_{j}}$. It follows then by ˜3.13(1) that $x^{\ell}-s(\gamma_{0})\prod_{-i\leq j\leq i}a_{j}^{h_{j}}$ is the minimal polynomial of $a_{i+1}$ over $\operatorname{VF}(A\langle a_{-i},\ldots,a_{i}\rangle)$, and $x^{\ell}-f(s(\gamma_{0}))\prod_{-i\leq j\leq i}b_{j}^{h_{j}}$ is the minimal polynomial of $b_{i+1}$ over $\operatorname{VF}(B\langle b_{-i},\ldots,b_{i}\rangle)$. Because we have $f_{i}(\prod_{-i\leq j\leq i}a_{j}^{h_{j}})=\prod_{-i\leq j\leq i}b_{j}^{h_{j}}$, we get that $f_{i}$ extends to a map
@@
If $\mathcal{L}=\mathcal{L}_{s,\iota,\sigma}$, the proof is analogous to the proof of the case $\mathcal{L}_{s,\iota,\sigma}$ in Lemma˜3.24 with the roles of the sorts $\Gamma$ and $k$ interchanged, the role of $\iota$ now played by $s$, and the applications of [Kesting, Proposition 1.7] replaced by applications of [Kesting, Proposition 2.1]. ∎
@@
Let $E,M,N$ be substructures of $\mathcal{L}$-valued difference fields of residual characteristic zero, with $M$ and $N$ extending $E$ and such that:
@@
Since $k(M)$ embeds into $k(N)$ over $k(E)$, there is an existentially closed (in particular, $\sigma$-henselian by Corollary˜3.11) and quantifier-free $\lvert M\rvert^{+}-$saturated extension $U$ of $M$ such that the restrictions to the algebraic closure of $k(E)$ of the automorphisms of $k(U)$ and $k(N)$ are isomorphic. By [CHr, Theorem 1.3] $\mathrm{id}_{E}$ is then a $k$-elementary map between substructures of $U$ and $N$.
@@
Observe that the proof of [DO, Corollary 5.10] [^14] For the reader’s convenience, we point out that most of the proofs in [DO, Section 5], which culminates in the corollary of interest, are omitted as they heavily rely on similar material presented in [AvD, Section 5]. only uses saturation to realise sets of quantifier-free formulas. Namely, the only step where saturation is used is in the proof of [AvD, Lemma 5.9], in saying that, in our notation, certain pseudo-Cauchy sequences from $B_{1}$ have pseudolimits (see for example [DO, Section 2] for the definitions of pseudo-Cauchy sequence, or pc-sequence, and of pseudolimit). To conclude, observe that if $\pi(x)$ is a set of quantifier-free formulas over $B_{1}$ stating that $x$ is a pseudolimit of a certain pc-sequence then, by definition of pc-sequence, $\pi(x)$ is finitely satisfiable in $B_{1}$, and a fortiori in $N$. ∎
As the extension $A_{1}\subseteq A_{2}$ is immediate, in order to conclude that $f_{2}$ is a $k$-elementary map it remains to be shown that, if $\mathcal{L}$ contains an angular component symbol, then $f_{2}$ commutes with $\operatorname{ac}$. This follows from the fact, because $a\in A_{2}$ is a pseudolimit of a pseudo-Cauchy sequence, there is $b\in A_{1}$ such that v(ba)>v(a)v(b-a)>v(a).
@@
An amalgamation problem of valued difference fields of residual characteristic zero with either an angular component or a cross-section, and possibly with a lift, has a solution if and only if the induced problem on the residue difference fields has a solution.